Bill sells is vintage 1974 Road Runner for $9500. He uses the money to invest in a 36 month CD that pays 4.12% interest compounded quartely. How much money will he receive when he cashes in his CD at the end of the 36 months?
Answer by stanbon(48558)

Bill sells is vintage 1974 Road Runner for $9500. He uses the money to invest in a 36 month CD that pays 4.12% interest compounded quartely. How much money will he receive when he cashes in his CD at the end of the 36 months?
-----------
A = P(1+r/n)^(nt)
Your Problem:
A is the future value
P is the pricipal or investment
r is the yearly interest rate
n is the number of compounding periods per year
t is the number of years
--------------------
A = 9500(1+0.0412/4)^(4*3)
A = 9500(1.1308)
A = $10743.06
==================
Cheers,
Stan H.